# supplementary

### v1

@@ 3.2. Statistical Inference Over Multiple Datasets
for all $i=0,\ldots,p$ and $j=1,\ldots,k$. In other words, if the variance structure of the spatial analysis model matches the data generation model and the conditional mean of $Z$ is a linear transformation of $X$, then $\widetilde{\beta}_{S}$ is always a more efficient estimator than $\widehat{\beta}_{NS}$ in terms of mean squared error over repeated realizations of $(X_{1},Y,Z)$. In multivariate analysis, this result depends both on the correct specification of the latent spatial variance structure and the correlation between outcomes. Paciorek (2010) argues that a classical spatial model does not adjust for confounding because it does not reduce bias compared to ordinary least squares. In our setting, this is true when the conditional mean of $Z$ is a linear transformation of $X$, a common assumption in research on spatial confounding from the data generation perspective (Thaden and Kneib, 2018; Dupont et al., 2022), but the benefit of the spatial analysis model is a potential reduction in the estimator’s variance, not bias.
The conditions for (13) hold for the data generation model discussed in (2) and (3) with $\delta=(\delta_{0},\delta_{1}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. So, the estimator $\widetilde{\beta}_{S}=\mathbb{E}[\beta_{S}\,|\,Y,M=A,R=P,X_{1}]$ from the spatial analysis model is preferable to $\widehat{\beta}_{NS}$. If $V_{\phi}^{-1}=W_{\phi}^{-1}$, the variance of $\widetilde{\beta}_{S}$ in the data generation model is equal to the posterior variance in the analysis model given the true data generation parameters,
$\operatorname{Var}(\operatorname{vec}(\widetilde{\beta}_{S})\,|\,X_{1})=\operatorname{Var}(\operatorname{vec}(\beta_{S})\,|\,Y,M=A,R=P,X_{1}),$ (14)
@@
where the left term is evaluated under the data generation model in (2) and the right term is evaluated under the spatial analysis model in (6) (see Appendix C for a proof). If $\widetilde{\beta}_{S}$ is unbiased (i.e. $\delta=0_{(p+1)\times k}$), credible intervals constructed using the conditional posterior will have coverage probability equal to the nominal coverage rate across realizations of $(Z,Y)$ with $X_{1}$ fixed. However, $\widetilde{\beta}_{S}$ cannot be calculated if $A$ and $P$ are not known. For fully Bayesian inference, one solution is to use the marginal posterior mean of $\beta$ under (6), which we denote as $\widehat{\beta}_{S}=\mathbb{E}[\beta_{S}\,|\,Y]$. A natural question is whether $\widehat{\beta}_{S}$ improves the estimation of $\beta$ under assumptions on the data-generating model similar to the case of $\widetilde{\beta}_{S}$. Because $\widehat{\beta}_{S}$ does not have a closed-form expression, we conduct simulation experiments to assess estimation performance and robustness of the analysis model under misspecified spatial structure.
@@ 4. Simulation Analysis of Spatial Estimator
We generate data sets $D^{(i)}=\left\{X_{1}^{(i)},Y^{(i)}\right\}$ using (2), where $X_{1}^{(i)}$ is a $58\times 1$ vector and $Y^{(i)}$ is a $58\times 3$ matrix of $k=3$ outcomes generated on a map of the $n=58$ California counties for $i=1,\ldots,300$. We set $\beta_{0}=(0,2,4)^{\mathrm{\scriptscriptstyle T}}$ and $\beta_{1}=(1,3,5)^{\mathrm{\scriptscriptstyle T}}$, $\delta_{0}=(0.4375,0.375,0.4375)^{\mathrm{\scriptscriptstyle T}}$, $\delta_{1}^{\mathrm{\scriptscriptstyle T}}=(0.125,0.25,0.125)^{\mathrm{\scriptscriptstyle T}}$, $\mu=0.5$, $\mu_{Z}=0.5$, $A=\begin{bmatrix}-0.5&-0.5&1.5\\ 1.25&-1.25&0\\ 1.5&1.5&1.5\end{bmatrix}$ such that $\Sigma=A^{\mathrm{\scriptscriptstyle T}}A=\begin{bmatrix}4.0625&0.9375&1.5\\ 0.9375&4.0625&1.5\\ 1.5&1.5&4.5\end{bmatrix}$, $P=\mbox{diag}(0.9,0.7,0.4)$, and $C=2$. To induce spatial dependence, we specify a conditionally autoregressive (CAR) structure, $V^{-1}_{\phi}=c(D_{W}-\alpha W)$, where $W$ is the adjacency matrix, $D_{W}$ is a diagonal matrix with each diagonal element $D_{W,ii}$ being the number of neighbors of county $i$, $\alpha$ is the spatial smoothing parameter, and $c$ is a scaling factor such that the geometric mean of the diagonal elements of $W_{\phi}^{-1}$ is equal to one. We set $\alpha=0.99$ and $c=0.8352$ to construct $V_{\phi}^{-1}$ using the map of the 58 Californian counties.
@@
For the spatial analysis model in (6), we employ the CAR prior through (7) with $W_{\phi}=V_{\phi}$, and refer to this model as the unconditioned CAR model. Following the suggestion of Jin et al. (2007), we set $v=3$ for all simulation runs and $\Sigma_{0}$ as a diagonal matrix with diagonal entries equal to $||(I_{n}-H)Y_{j}||^{2}/(n-p-1)$. Although the prior $\pi(\Sigma_{S})$ is dependent on the data, Jin et al. (2007) argues that this enables robust posterior inference. We fix $\lambda_{R}=0.01$ for $\pi(R)$, a vague prior on $R$ that attributes slightly higher prior density to smaller values of $r_{i}$. The spatial estimates $\widehat{\beta}_{S}^{(i)}$ and 95% highest posterior density credible intervals for each fixed effect are computed using 20,000 posterior samples of $\beta_{S}$ after 20,000 burn-in samples obtained through a Metropolis within Gibbs sampling scheme detailed in Appendix D with proposal settings $s_{1}=0.10$, $s_{2}=0.15$, and $s_{3}=0.25$. For the conditioned CAR model, we calculate $\widetilde{\beta}^{(i)}$ using (9) by setting $M$, $R$, and $W_{\phi}$ equal to the true simulation values of $A$, $P$, and $V_{\phi}$, respectively. We compute 95% conditional credible intervals as $\widetilde{\beta}_{ij}\pm z_{0.975}\sqrt{\operatorname{Var}(\beta_{S,ij}\,|\,Y,X_{1},A,P)}$ for $i=0,\ldots,p$ and $j=1,\ldots,k$. For the non-spatial model in (4), we set $\pi(\beta_{NS},\Sigma_{NS})\propto\mbox{IW}(\Sigma_{NS}\,|\,v,v\Sigma_{0})$ with the hyperparameters $v=3$ and $v\Sigma_{0}$ identical to the scale matrix for $\pi(\Sigma_{S})$. We calculate 95% highest posterior density credible intervals for each fixed effect using 10,000 exact posterior samples from $\pi(\beta_{NS},\Sigma_{NS}\,|\,Y,X_{1})$. To examine the effects of misspecifying the spatial prior, we evaluate the aforementioned estimates and credible intervals under a SAR spatial prior on $\gamma$ by setting $W_{\phi}^{-1}=c(I_{n}-\alpha\widetilde{W})(I_{n}-\alpha\widetilde{W})^{\mathrm{\scriptscriptstyle T}}$ where $\widetilde{W}$ is a row-normalization of $W$ such that $\widetilde{W}_{ij}=W_{ij}/\sum_{m=1}^{n}W_{im}$, $\alpha=0.99$ is the spatial smoothing parameter, and $c=220.1809$ is a scaling factor such that the geometric mean of the prior marginal variances is equal to one.
Frequentist Evaluation $\beta_{01}$ $\beta_{02}$ $\beta_{03}$ $\beta_{11}$ $\beta_{12}$ $\beta_{13}$
$\mbox{MSE}(\mathbb{E}[\beta_{NS}\,|\,Y,X_{1}])$ 1.513 1.168 1.517 0.076 0.111 0.095
$\mbox{MSE}(\mathbb{E}[\beta_{S}\,|\,Y,X_{1}])$, CAR Prior 1.435 1.065 1.453 0.050 0.090 0.054
$\mbox{MSE}(\mathbb{E}[\beta_{S}\,|\,Y,X_{1}])$, SAR Prior 1.507 1.131 1.523 0.056 0.095 0.062
$\mbox{MSE}(\mathbb{E}[\beta_{S}\,|\,Y,X_{1},A,P])$, CAR Prior 1.401 1.062 1.458 0.049 0.090 0.050
$\mbox{MSE}(\mathbb{E}[\beta_{S}\,|\,Y,X_{1},A,P])$, SAR Prior 1.516 1.167 1.687 0.054 0.095 0.056
$\mbox{Coverage}(\beta_{NS})$ 0.373 0.433 0.423 0.723 0.593 0.667
$\mbox{Coverage}(\beta_{S})$, CAR Prior 0.867 0.923 0.873 0.917 0.763 0.887
$\mbox{Coverage}(\beta_{S})$, SAR Prior 0.983 0.993 0.963 0.867 0.683 0.830
$\mbox{Coverage}(\beta_{S}\,|\,A,P)$, CAR Prior 0.907 0.950 0.940 0.887 0.713 0.887
$\mbox{Coverage}(\beta_{S}\,|\,A,P)$, SAR Prior 0.980 0.993 0.973 0.767 0.523 0.690
$\mathbb{E}[\operatorname{Var}(\beta_{NS}\,|\,Y,X_{1})]$ 0.109 0.108 0.108 0.027 0.026 0.027
$\mathbb{E}[\operatorname{Var}(\beta_{S}\,|\,Y,X_{1})]$, CAR Prior 1.006 0.963 0.937 0.039 0.038 0.037
$\mathbb{E}[\operatorname{Var}(\beta_{S}\,|\,Y,X_{1})]$, SAR Prior 2.082 2.100 2.081 0.032 0.032 0.032
$\mathbb{E}[\operatorname{Var}(\beta_{S}\,|\,Y,X_{1},A,P)]$, CAR Prior 1.054 1.054 1.353 0.034 0.034 0.034
$\mathbb{E}[\operatorname{Var}(\beta_{S}\,|\,Y,X_{1},A,P)]$, SAR Prior 1.983 1.983 2.569 0.020 0.020 0.017
@@
Table 1. Frequentist evaluation of point estimates and 95% credible intervals from non-spatial and spatial CAR models across 300 datasets generated under a CAR spatial data generation model on a map of Californian counties.
[…]
@@
For the SAR analysis, the efficiency result in (13) is not guaranteed as the analysis spatial structure $W_{\phi}^{-1}$ does not match $V_{\phi}^{-1}$ in the true data generating model. Nonetheless, the unconditioned spatial model obtains lower mean squared errors and credible interval coverage rates closer to the nominal level compared to the non-spatial model (Table 1). Therefore, a confounded and misspecified spatial effect may still be preferable to none for estimation, albeit at the cost of large posterior variance inflation for the intercept terms. In practice, we recommend fitting multiple spatial models and using an information criteria such as DIC (Spiegelhalter et al., 2002) or WAIC (Watanabe, 2010) to perform model selection.
@@
We assess model fit by computing the Kullback-Leiber divergence $D_{KL}(p(Y_{0})||q(Y_{0}))$ for each dataset and analysis model. Here, $Y_{0}$ is an observation from a new realization of the underlying stochastic process conditional on $X_{1}$, $p(Y_{0})$ is the data generation density, and $q(Y_{0})$ is the predictive density under a analysis model. Marginalizing out $Z$ in the data generation model from (2) yields $p(Y_{0})=\mbox{N}_{nk}(\operatorname{vec}(Y_{0})\,|\,X(\beta+\delta),A^{\mathrm{\scriptscriptstyle T}}(I_{k}-P)A\otimes I_{n}+A^{\mathrm{\scriptscriptstyle T}}PA\otimes V_{\phi})$. In the non-spatial model, $q(Y_{0})=\mbox{N}_{nk}(\operatorname{vec}(Y_{0})\,|\,\operatorname{vec}(X\beta_{NS}),\Sigma_{NS}\otimes I_{n})$. For the conditioned spatial model, $q(Y_{0})=\mbox{N}_{nk}(\operatorname{vec}(Y_{0})\,|\,\operatorname{vec}(X\beta_{S}),A^{\mathrm{\scriptscriptstyle T}}(I_{k}-P)A\otimes I_{n}+A^{\mathrm{\scriptscriptstyle T}}PA\otimes W_{\phi})$, and for the unconditioned spatial model, $q(Y_{0}\,|\,X_{1})=\mbox{N}_{nk}(\operatorname{vec}(Y_{0})\,|\,\operatorname{vec}(X\beta_{S}),M^{\mathrm{\scriptscriptstyle T}}(I_{k}-R)M\otimes I_{n}+M^{\mathrm{\scriptscriptstyle T}}RM\otimes W_{\phi})$. We use the aforementioned posterior samples for the non-spatial and unconditioned spatial model along with 20,000 exact samples of $\beta_{S}$ in the conditioned spatial model to obtain samples from the posterior distribution of $D_{KL}(p(Y_{0})||q(Y_{0}))$ for each dataset and analysis model. The density plot of the posterior means in Figure 1 demonstrates that the unconditioned spatial models generally provide a better fit than the non-spatial model. Therefore, the increased posterior variance in the spatial models is a more accurate representation of the true variability in the data generation model. The unconditioned SAR model outperforms the conditioned SAR model, adapting to the misspecified spatial structure.

### v2

@@ 3.3. Estimation Over Multiple Datasets
for all $i=1,\ldots,p+1$ and $j=1,\ldots,k$. In other words, if the variance structure of the spatial analysis model matches the data generation model and the conditional mean of $Z$ is a linear transformation of $X$, then $\widetilde{B}_{S}$ is a more efficient estimator than $\widehat{B}_{NS}$ in terms of mean squared error over repeated realizations of $(X_{1},Y,Z)$. This result depends on correct specification of both the latent spatial variance structure and the correlation between outcomes.
Uncertainty quantification in the spatial BYM2 model matches the variance of the point estimator under the data generation process. If $V_{\phi}^{-1}=W_{\phi}^{-1}$, the variance of $\widetilde{B}_{S}$ in the data generation model conditional on $X_{1}$ is equal to the posterior variance in the analysis model,
$\operatorname{Var}_{DG}(\operatorname{vec}(\widetilde{B}_{S})\,|\,X_{1})=\operatorname{Var}(\operatorname{vec}(B_{S})\,|\,Y,M=A,R=P,X_{1}),$ (12)
where the left-hand variance term is evaluated under the data generation model in (3) and the right-hand term is evaluated under the spatial analysis model in (6) (see Section LABEL:append:spatial_variance for a proof). Credible intervals for the unconditional estimand $F=B+D$ constructed using the conditional posterior will have coverage probability equal to the nominal coverage rate across realizations of $(Z,Y)$ with $X_{1}$ fixed. Combining (11) and (12) implies that approaches such as restricted spatial regression methods that preserve the non-spatial estimates but reduce the posterior variance result in credible intervals that do not have a valid uncertainty interpretation under the data generation process. This aligns with previous results by Zimmerman and Ver Hoef (2022); Khan and Calder (2022) that show restricted spatial regression methods tend to underestimate the true uncertainty in estimation of fixed effects.
@@
However, $\widetilde{B}_{S}$ cannot be calculated if $A$ and $P$ are not known. For fully Bayesian inference, one solution is to use the marginal posterior mean of $B$ under (6), which we denote as $\widehat{B}_{S}=\mathbb{E}[B_{S}\,|\,Y,X]$. Because $\widehat{B}_{S}$ does not have a closed-form expression, we conduct simulation experiments to assess estimation performance and robustness of the analysis model when between-outcome dependence is unknown and spatial structure is misspecified.
@@ 4. Simulation Analysis of Spatial Estimator
We generate data sets $D^{(i)}=\left\{X_{1}^{(i)},Y^{(i)}\right\}$ using (3), where $X_{1}^{(i)}$ is a $58\times 1$ vector and $Y^{(i)}$ is a $58\times 2$ matrix of $k=2$ outcomes generated on a map of the $n=58$ California counties for $i=1,\ldots,300$. We set $\beta_{0}=(0,2)^{\mathrm{\scriptscriptstyle T}}$, $B_{1}=(1,3)^{\mathrm{\scriptscriptstyle T}}$, $\delta_{0}=(0.5,0.5)^{\mathrm{\scriptscriptstyle T}}$, $D_{1}=(0.3,0.3)^{\mathrm{\scriptscriptstyle T}}$, $\mu=0.5$, $A=\begin{bmatrix}1&0.5\\ 0.5&1.0\\ \end{bmatrix}$ such that $\Sigma=A^{\mathrm{\scriptscriptstyle T}}A=\begin{bmatrix}1.5&1.0\\ 1.0&1.5\\ \end{bmatrix}$, $P=\mbox{diag}(0.9,0.7)$, and $C=2$. We generate the data using a conditionally autoregressive (CAR) structure, $V^{-1}_{\phi}=c(D_{W}-\alpha W)$, where $W$ is the adjacency matrix, $D_{W}$ is a diagonal matrix with each diagonal element $D_{W,ii}$ being the number of neighbors of county $i$, $\alpha$ is the spatial smoothing parameter, and $c$ is a scaling factor such that the geometric mean of the diagonal elements of $W_{\phi}^{-1}$ is equal to one. We set $\alpha=0.99$ and $c=0.8352$ to construct $V_{\phi}^{-1}$ using the Californian county map.
@@
For the spatial analysis model in (6), we employ the CAR prior through (7) with $W_{\phi}=V_{\phi}$, and refer to this model as the unconditioned CAR model. Following Jin et al. (2007), we set $v=2$ for all simulation runs and $\Sigma_{0}$ as a diagonal matrix with diagonal entries equal to $(I_{n}-H)Y_{j}||^{2}/(n-p-1)$. Although the prior $\pi(\Sigma_{S})$ is dependent on the data, Jin et al. (2007) argues that this enables robust posterior inference. We fix $\lambda_{R}=0.01$ for $\pi(R)$, a vague prior on $R$ that attributes slightly higher prior density to smaller values of $r_{i}$. The spatial estimates $\widehat{B}_{S}^{(i)}$ and 95% highest posterior density credible intervals for each fixed effect are computed using 20,000 posterior samples after 20,000 burn-in samples obtained through a Metropolis within Gibbs sampling scheme detailed in Section LABEL:append:spatial_sampling_scheme with proposal settings $s_{1}=0.10$, $s_{2}=0.15$, and $s_{3}=0.25$. For the conditioned CAR model, we calculate $\widetilde{B}^{(i)}$ using (8) by setting $M$, $R$, and $W_{\phi}$ equal to the true values of $A$, $P$, and $V_{\phi}$, respectively.
We compute 95% conditional credible intervals for the elements of $F=B+D$ as $\widetilde{B}_{ij}\pm z_{0.975}\sqrt{\operatorname{Var}(B_{S,ij}\,|\,Y,M=A,R=P,X_{1})}$ for $i=0,\ldots,p$ and $j=1,\ldots,k$. For the non-spatial model in (4), we set $\pi(B_{NS},\Sigma_{NS})\propto\mbox{IW}(\Sigma_{NS}\,|\,v,v\Sigma_{0})$ with the hyperparameters $v=2$ and $v\Sigma_{0}$ identical to the scale matrix for $\pi(\Sigma_{S})$. We calculate 95% highest posterior density credible intervals for each fixed effect using 10,000 exact posterior samples from $\pi(B_{NS},\Sigma_{NS}\,|\,Y,X_{1})$. To examine the effects of misspecifying the spatial prior, we evaluate the aforementioned estimates and credible intervals under a SAR spatial prior on $G$ in (6) by setting $W_{\phi}^{-1}=c(I_{n}-\alpha\widetilde{W})(I_{n}-\alpha\widetilde{W})^{\mathrm{\scriptscriptstyle T}}$ where $\widetilde{W}$ is a row-normalization of $W$ such that $\widetilde{W}_{ij}=W_{ij}/\sum_{m=1}^{n}W_{im}$, $\alpha=0.99$ is the spatial smoothing parameter, and $c=220.1809$ is a scaling factor such that the geometric mean of the prior marginal variances is equal to one.
Frequentist Evaluation Model $F_{11}$ $F_{12}$ $F_{21}$ $F_{22}$
Mean-squared Error $B_{NS}\,|\,Y,X_{1}$ 0.642 0.592 0.112 0.109
$B_{S}\,|\,Y,X_{1},A,P$; CAR Prior 0.634 0.588 0.098 0.099
$B_{S}\,|\,Y,X_{1},A,P$; SAR Prior 0.635 0.598 0.103 0.104
$B_{S}\,|\,Y,X_{1}$; CAR Prior 0.635 0.585 0.096 0.098
$B_{S}\,|\,Y,X_{1}$; SAR Prior 0.674 0.612 0.101 0.103
Coverage of $F=B+D$ $B_{NS}\,|\,Y,X_{1}$ 0.357 0.370 0.657 0.713
$B_{S}\,|\,Y,X_{1},A,P$; CAR Prior 0.800 0.863 0.917 0.923
$B_{S}\,|\,Y,X_{1},A,P$; SAR Prior 0.930 0.967 0.837 0.847
$B_{S}\,|\,Y,X_{1}$; CAR Prior 0.973 0.970 0.950 0.933
$B_{S}\,|\,Y,X_{1}$; SAR Prior 0.990 0.997 0.613 0.690
Average Posterior Variance $B_{NS}\,|\,Y,X_{1}$ 0.026 0.027 0.007 0.007
$B_{S}\,|\,Y,X_{1},A,P$; CAR Prior 0.185 0.221 0.008 0.009
$B_{S}\,|\,Y,X_{1},A,P$; SAR Prior 0.417 0.476 0.007 0.008
$B_{S}\,|\,Y,X_{1}$; CAR Prior 0.480 0.420 0.008 0.009
$B_{S}\,|\,Y,X_{1}$; SAR Prior 0.927 0.808 0.002 0.004
@@
Table 3. Frequentist evaluation of point estimates and 95% credible intervals for $F=B+D$ from non-spatial/spatial CAR and SAR models across 300 datasets generated under a CAR spatial data generation model on a map of Californian counties.
The frequentist evaluation of the posterior means and 95% credible intervals for each analysis model in Table 3 provides simulation evidence for the claim of (11), as the conditional spatial estimator $\widetilde{B}_{S}$ is superior to $\widehat{B}_{NS}$ in terms of mean squared error for every coefficient despite the presence of the spatial confounder $Z$ in the data generation model. The spatial models show inflated posterior variances for the intercept terms compared to the non-spatial model. The conditioned and unconditioned CAR model have similar mean-squared errors, while the credible intervals of the conditioned CAR model exhibit modest undercoverage. The unconditioned spatial CAR prior provides frequentist coverage closest to the nominal credible level, while the non-spatial model does not yield a coverage rate above 75% for any coefficient. These results illustrate that the multivariate spatial analysis model in (6) can improve estimation compared to the non-spatial analysis model in (4) despite the effects of spatial confounding and uncertainty about the variance parameters $A$ and $P$. Increased posterior variance in the spatial model compared to the non-spatial model is not a inherent flaw, but rather a reason to distrust the non-spatial model which suffers from undercoverage.
@@
For the SAR analysis, the efficiency result in (11) is not guaranteed as the analysis spatial structure $W_{\phi}^{-1}$ does not match $V_{\phi}^{-1}$ in the true data generating model. Nonetheless, the unconditioned spatial model obtains lower mean squared errors and credible interval coverage rates closer to the nominal level compared to the non-spatial model (Table 3). Therefore, a confounded and misspecified spatial effect may still be preferable to none for estimation, albeit at the cost of large posterior variance inflation for the intercept terms and potentially inaccurate uncertainty quantification. In practice, we recommend fitting multiple spatial models and using an information criteria such as DIC (Spiegelhalter et al., 2002) or WAIC (Watanabe, 2010) to perform model selection.
@@
We assess model fit by computing the Kullback-Leiber divergence $D_{KL}(p(Y_{0})||q(Y_{0}))$ for each dataset and analysis model. Here, $Y_{0}$ is an observation from a new realization of the underlying stochastic process conditional on $X_{1}$, $p(Y_{0})$ is the data generation density, and $q(Y_{0})$ is the predictive density under a analysis model. Marginalizing out $Z$ in the data generation model from (3) yields $p(Y_{0})=\mbox{N}_{nk}(\operatorname{vec}(Y_{0})\,|\,X(B+D),A^{\mathrm{\scriptscriptstyle T}}(I_{k}-P)A\otimes I_{n}+A^{\mathrm{\scriptscriptstyle T}}PA\otimes V_{\phi})$. In the non-spatial model, $q(Y_{0})=\mbox{N}_{nk}(\operatorname{vec}(Y_{0})\,|\,\operatorname{vec}(XB_{NS}),\Sigma_{NS}\otimes I_{n})$. For the conditioned spatial model, $q(Y_{0})=\mbox{N}_{nk}(\operatorname{vec}(Y_{0})\,|\,\operatorname{vec}(XB_{S}),A^{\mathrm{\scriptscriptstyle T}}(I_{k}-P)A\otimes I_{n}+A^{\mathrm{\scriptscriptstyle T}}PA\otimes W_{\phi})$, and for the unconditioned spatial model, $q(Y_{0}\,|\,X_{1})=\mbox{N}_{nk}(\operatorname{vec}(Y_{0})\,|\,\operatorname{vec}(XB_{S}),M^{\mathrm{\scriptscriptstyle T}}(I_{k}-R)M\otimes I_{n}+M^{\mathrm{\scriptscriptstyle T}}RM\otimes W_{\phi})$. We use the aforementioned posterior samples for the non-spatial and unconditioned spatial model along with 20,000 exact samples of $B_{S}$ in the conditioned spatial model to obtain samples from the posterior distribution of $D_{KL}(p(Y_{0})||q(Y_{0}))$ for each dataset and analysis model. The density plot of the posterior means in Figure 1 demonstrates that the unconditioned spatial models generally provide a better fit than the non-spatial model. Therefore, the increased posterior variance in the spatial models is a more accurate representation of the true variability in the data generation model. The unconditioned SAR model outperforms the conditioned SAR model, adapting to the misspecified spatial structure.